OpenStudy (anonymous):

Can someone help me with this equation? I'm trying but it's not coming out right.

OpenStudy (anonymous):

Find an exact solution for Then, find the approximate solution.

OpenStudy (anonymous):

start by "cross multiplying" to get \[2(\sqrt{5}-1)=\sqrt{5}x\]

OpenStudy (anonymous):

\[x=\frac{2(\sqrt{5}-1)}{\sqrt{5}}\]

OpenStudy (anonymous):

\[2\sqrt{5}-2=\sqrt{5}x\] \[x=\frac{2\sqrt{5}-2}{\sqrt{5}}\]

OpenStudy (anonymous):

maybe you need to "rationalize the denominator" as well

OpenStudy (anonymous):

multiply and divide. You just have to keep track of things. It's better when you put parentheses on things when you do this.

OpenStudy (anonymous):

okay, i'll try again..

OpenStudy (anonymous):

I'm still getting an answer thats not on my choice list.. :/

OpenStudy (anonymous):

what are ur choices????

OpenStudy (anonymous):

a. 10-2 sqrt5/5; 1.1 b. 2-2sqrt5; -2.5 c. -2; 2 d. 2sqrt5-2/sqrt5; 1.1

OpenStudy (anonymous):

ya its d

OpenStudy (anonymous):

can you explain how you did and it got that answer please?

OpenStudy (anonymous):

see the first part of d has already been solved by satellite 73

OpenStudy (anonymous):

for 1.1,you can simply put the value of sqrt(5) and solve....

OpenStudy (anonymous):

Oh i see.. lol ty